OpenStudy (anonymous):

Particle 1 has a velocity v1 = 2 m/s and a mass m1 = 2 kg . This particle collides with particle 2 of mass m2 = 6 kg , which is at rest, and then the two particles stick together. The speed of the conglomerate particle is 1. 2 m/s 2. 1 m/s 3. 0.5 m/s 4. 0.75 m/s 5. 1.5 m/s

OpenStudy (anonymous):

The ratio of the final kinetic energy to the initial kinetic energy is 1. 0.25 2. 0.1 3. 0.8 4. 1 5. 0.5

OpenStudy (primeralph):

m1v1+m2v2 = (m1+m2)V Find V.

OpenStudy (amistre64):

reminds me of a weighted average .....

OpenStudy (anonymous):

so it's (2*2)+(6*0)=(2+6)V

OpenStudy (primeralph):

Yes.

OpenStudy (primeralph):

@amistre64 It IS a weighted average. It's just a form of the expected velocity. It's interesting how topics branch in and out.

OpenStudy (amistre64):

agreed :)

OpenStudy (anonymous):

So the answer would be .5m/s?

OpenStudy (primeralph):

If you did it right, yes. Remember to use the final mass for the KE.

OpenStudy (anonymous):

And to find kinetic energy I use .5mv^2? But there are two masses

OpenStudy (primeralph):

Find KE 1 and KE 2 and then KE final.

OpenStudy (anonymous):

Oh ok but I add the first two?

OpenStudy (primeralph):

Yes.

OpenStudy (primeralph):

The second mass has 0 KE since it is at rest.

OpenStudy (anonymous):

Thank you :)

OpenStudy (primeralph):

You're welcome.